OpenStudy (anonymous):

can you please show me the steps to solving (x-3)(x+5)=9

OpenStudy (anonymous):

Multiply (x - 3)(x + 5) = x^2 + 2x - 15 = 9 Then it must = 0 so x^2 + 2x - 24 = 0 Now you can factor (x + 6)(x - 4) = 0 x = -6, 4